# Two-Time Relativistic Bohmian Model of Quantum Mechanics

### v7

@@ Abstract
Two-Time relativistic Bohmian Model (TTBM) is a theory in which the apparently paradoxical aspects of Quantum Mechanics are the effect of the existence of an extra unobservable time dimension. The kind of matter ubiquity suggested by Feynman’s sum over history formulation has already been confirmed by violations of the Leggett-Garg inequality and quantum teleportation. The hypothesis that matter is capable of motion with respect to an extra time (thus being instantaneous with respect to usual time) is capable of restoring determinism, explaining the Zitterbewegung without evoking antimatter. The model also predicts a relativistic correction of the uncertainty principle.
Here the model is first summarized (definition, salient properties, and empiricism) and after applied to a generic spherical atomic orbit, obtaining electron oscillations in the new time dimension, $\tau$, which demonstrate the static nature of the orbitals. Something very similar happens in the case of a particle in a box, where $\tau$-oscillations cause the particle to spread out at steady states. Some astrophysical and about the spin speculations follow. Finally, it is discussed how the model fits into the fundamental problem of the definition of time in Quantum Mechanics.
@@ I Introduction
Several experiments on violations of the Leggett-Garg inequality [1, 2, 3, 4] and quantum teleportation [5, 6, 7, 8] have long demonstrated that the kind of ubiquity envisioned by Feynman’s interpretation of the sum of histories [9, 10, 11] - a formulation equivalent to the standard one - is real. All the apparent paradoxes of Quantum Mechanics are rooted in it. However, to restore determinism, all that’s needed is to introduce an extra temporal dimension: a measure that, though shocking at first glance, is physically equivalent to the introduction of additional spatial dimensions, already contemplated by several modern theories.
The TTBM theory, introduced recently [12, 13], is linked to the attempts at relativistic generalization of the de Broglie-Bohm mechanics [14, 15, 16, 17, 18, 19]. One of the crucial points in this effort is the definition of simultaneity [20]. Until a couple of years ago, the most successful strategy for this purpose was to introduce models which, while respecting Lorentz invariance, must admit superluminal speeds and complex space-time structures [21, 22, 23, 24, 25, 26, 27, 28]. The approach of TTBM is more elementary: an additional time $\tau$, independent of the usual time $t$ is introduced. Motions in $\tau$ are instantaneous with respect to t time and, while not directly observable, are supposed to generate the quantum uncertainties of the observables. They are caused by a Bohmian-type Quantum Potential and find the basic example in the famous Zitterbewegung motion [29, 30], of which a more general law, free from the call of antimatter, is achieved. Beyond the empiricism of standard Quantum Mechanics, the model foresees a relativistic anisotropic modification of the uncertainty principle, which could be tested in high-energy accelerators.
@@ II TTBM definition and properties
Two-time physics has been considered in various contexts. An additional temporal parameter is considered, e.g., in [31, 32, 33, 34, 35, 36, 37] and, with a different approach, in [38, 39, 40, 41] in order to more accurately explain different quantum properties or obtain supersymmetries. But in all these theories, the expectation value of the new time is not independent from the old one.
@@
Eq. (2) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to these surprising facts: electrons spread in their atomic orbital, quantum diffraction, entanglement, violations of the Leggett-Garg inequality, teleportation, etc. Really, it should be like this in any motion: according to Feynman’s sum over histories interpretation, equivalent to the standard one, the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled all the possible paths from A to B at once, also interacting with every possible virtual particle of the vacuum [10, 11]. In ubiquity lie all the puzzling aspects of Quantum Mechanics but the mere consideration of extra time is capable of explaining it deterministically.
@@
So in TTBM, the Bohm’s original motion equation [16]: $\frac{{\rm d}\vec{p}}{{\rm d}t}=-\vec{\nabla}V-\vec{\nabla}V_{Q}$ does not hold: the motion is instead described by eqs. (11) and, for each spatial direction $\hat{s}$, (13). These corpuscular equations are complemented by the Klein-Gordon wave equation.
@@
Apparently, the model adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we shall see in the next sections, the motions in the new time variable are normally negligible for macroscopic objects. Conversely, for corpuscles, it happens that even assuming to know the Quantum Potential with a good approximation, due to the independence of the two times and the omnidirectionality of the $\tau$-motions, the function $\vec{r}(t,\tau)$ does not represent the commonly understood trajectory, but rather a three-dimensional object of $t$-spacetime: while $t$ remains constant, we will find that the particle can $\tau$-oscillate an arbitrary number of times in any direction. However, this is an indeterminism linked to the nonexistence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model, we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them, but the wave point of view is necessary due to the described constitutional insufficiency of the corpuscular description.
@@ III Free particle
Finally, it’s time to discuss an important assumption of the theory: that oscillatory motions similar to those found also occur in the case of an external potential $V$. The reasons are generally valid: the equations imply that $V$ modifies only indirectly, and usually slightly, the Quantum Potential. Exceptions could arise for potentials comparable to $m\gamma_{c}c^{2}$, therefore at the quark level (in which case there would be significant variations in quantum uncertainties). However later we will explicitly find the oscillatory $\tau$-motions of electrons in an atomic orbit.
@@ IV TTBM empiricism
Finally, we heuristically describe the modalities of Entanglement of multiple particles. According to TTBM, however distant the particles may be from each other and how weak their entanglement, all the particles must share an entire volume of connection. That is, each particle, through its $\tau$-motions, is spread out in this volume and can be observed anywhere within it (this does not mean that in certain areas this probability cannot be arbitrarily small). Measurement of a particle’s property will cause a phase change in a subset of wavefunctions. This will result in $\tau$-interferences with other wavefunctions of the system, capable of transmitting information - istantly with respect to the usual time - to any other particle.
@@ V Atomic orbitals
The standard approach, which considers stationarity ($\frac{\partial R}{\partial t}=0$) and $V_{Q}=0$, is capable of determining the quantized orbits, as in a hydrogen-like atom; but only inside the new model it possible to demonstrate, finally, the static nature of the orbitals of any atom. We will limit ourselves to analyzing spherical orbitals, so getting, in a polar reference system with center in the nucleus, just a dependence $R=R(r,\theta,t,\tau)$.

### v8

@@ Abstract
Two-Time relativistic Bohmian Model (TTBM) is a theory in which the apparently paradoxical aspects of Quantum Mechanics are the effect of the existence of an extra unobservable time dimension. The hypothesis that matter is capable of motion with respect to an additional independent time (thus being instantaneous with respect to usual time) is capable of restoring determinism, explaining the Zitterbewegung without evoking antimatter. The model also predicts a relativistic correction of the uncertainty principle.
Here the model is first summarized (definition, salient properties and empiricism) and after applied to a generic spherical atomic orbit, obtaining electron oscillations in the new time dimension, $\tau$, which demonstrate the static nature of the orbitals. Something very similar happens in the case of a particle in a box, where $\tau$-oscillations cause the particle to spread out at steady states. Some astrophysical and about the spin speculations follow. Finally, it is discussed how the model fits into the fundamental problem of the definition of time in Quantum Mechanics.
@@ I Introduction
In Feynman’s sum of histories interpretation [1, 2, 3] - a formulation equivalent to the standard one - all the apparent paradoxes of Quantum Mechanics converge into a single surprising aspect: the ubiquity of matter. Yet, to resolve it and restore determinism, all that’s needed is to introduce an extra temporal dimension: a measure that, though shocking at first glance, is physically equivalent to the introduction of additional spatial dimensions, already contemplated by several modern theories. This is achieved in the recently introduced TTBM theory [4, 5], linked to the attempts at relativistic generalization of the de Broglie-Bohm mechanics [6, 7, 8, 9, 10, 11]. The crucial point in this effort is the definition of simultaneity [12]. Until a couple of years ago, the most successful strategy for this purpose was to introduce models which, while respecting Lorentz invariance, must admit superluminal speeds and complex space-time structures [13, 14, 15, 16, 17, 18, 19, 20]. The approach of TTBM is more elementary: an additional time $\tau$, independent of the usual time $t$ is introduced. Motions in $\tau$ are instantaneous with respect to t time and, while not directly observable, are supposed to generate the quantum uncertainties of the observables. They are caused by a Bohmian-type Quantum Potential and find the basic example in the famous Zitterbewegung motion [21, 22], of which a more general law, free from the call of antimatter, is achieved. Beyond the empiricism of standard Quantum Mechanics, the model foresees a relativistic anisotropic modification of the uncertainty principle, which could be tested in high-energy accelerators.
@@ II TTBM definition and properties
Two-time physics has been considered in various contexts. An additional temporal parameter is considered, e.g., in [23, 24, 25, 26, 27, 28, 29] and, with a different approach, in [30, 31, 32, 33, 34, 35] in order to more accurately explain different quantum properties or obtain supersymmetries. But in all these theories, the expectation value of the new time is not independent from the old one.
@@
Eq. (2) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to these surprising facts as the spreading of the electrons in their atomic orbitals and the quantum diffractions. Really, it should be like this in any motion: according to Feynman’s sum over histories interpretation, equivalent to the standard one, the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled all the possible paths from A to B at once, also interacting with every possible virtual particle of the vacuum [2, 3]. In ubiquity lie all the puzzling aspects of Quantum Mechanics but the mere consideration of extra time is capable of explaining it deterministically.
@@
So in TTBM, the Bohm’s original motion equation [8]: $\frac{{\rm d}\vec{p}}{{\rm d}t}=-\vec{\nabla}V-\vec{\nabla}V_{Q}$ does not hold; the motion is instead described by eqs. (11) and, for each spatial direction $\hat{s}$, (13). These corpuscular equations are complemented by the Klein-Gordon wave equation.
@@
Apparently, the model adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we shall see in the next sections, the motions in the new time variable are normally negligible for macroscopic objects. Conversely, for corpuscles, it happens that even assuming to know the Quantum Potential with a good approximation, due to the independence of the two times and the omnidirectionality of the $\tau$-motions, the function $\vec{r}(t,\tau)$ does not represent the commonly understood trajectory, but rather a three-dimensional object of $t$-spacetime: while $t$ remains constant, we will find that the particle can $\tau$-oscillate an arbitrary number of times in any direction. However, this is an indeterminism linked to the nonexistence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model, we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them: simply, the wave point of view is necessary due to the described constitutional insufficiency of the corpuscular description.
@@ III Free particle
Finally, it’s time to discuss an important assumption of the theory: that oscillatory motions similar to those found also occur in the case of an external potential $V$. The reasons are generally valid: the equations imply that $V$ modifies only indirectly, and usually slightly, the Quantum Potential. Exceptions could arise for potentials comparable to $m\gamma_{c}c^{2}$, therefore at the quark level (in which case there would be significant variations in quantum uncertainties). However later we will explicitly find the oscillatory $\tau$-motions of electrons in a circular atomic orbit.
@@ IV TTBM empiricism
Finally, we heuristically describe the modalities of Entanglement of multiple particles. According to TTBM, however distant the particles may be from each other and how weak their entanglement, all the particles must share an entire volume of connection. That is, each particle, through its $\tau$-motions, is spread out in this volume and can be observed anywhere within it (this does not mean that in certain areas this probability cannot be arbitrarily small). Measurement of a particle’s property will cause a phase change in a subset of wavefunctions. This will result in $\tau$-interferences with other wavefunctions of the system, capable of transmitting information - istantly with respect to the usual time - to any other particle. Recall that this explicit non-locality has been confirmed by experiments [38, 39, 40].
@@ V Atomic orbitals
The standard approach, which considers stationarity ($\frac{\partial R}{\partial t}=0$) and $V_{Q}=0$, is capable of determining the quantized orbits, as in a hydrogen-like atom; but only inside the new model it possible to demonstrate, finally, the static nature of the orbitals of any atom. We will limit ourselves to analyzing the case of a circular motion, so getting, in a polar reference system with center in the nucleus, just a dependence $R=R(r,\theta,t,\tau)$.